# Prompting with Divide-and-Conquer Program Makes Large Language Models Discerning to Hallucination and Deception

### v4

@@ 4.1 Fixed-Depth Transformer can not solve 2-color Binary Subtree Isomorphism (2-BSI)
2-color Binary Subtree Isomorphism problem is that, given a pattern 2-color1112-color means that each node in the tree can be of two colors. We can understand it as that each node has a 0-1 label binary tree $t_{p}$ and a base 2-color binary tree $t_{b}$, a solver is required to judge whether the pattern tree is isomorphic to a sub-tree of $t_{b}$
@@
Assume that $TC^{0}\not=NC^{1}$. For any depth $L$, any polynomial $Q$, there exists a size $n$ of pattern tree such that there exist no log-precision Transformer with a depth of $L$, hidden dimension $d<Q(n)$, and fixed prompt $p$ that can directly output the solution (Yes or No) of the 2-color Binary Subtree Isomorphism problem (2-BSI).
@@
The core point of the proof is that, based on (Merrill & Sabharwal, 2023a), we know that the expressive power of log-precision Transformer is in $TC^{0}$. As proved by (Jenner et al., 2003), any $L$ problem ($NC^{1}\subseteq L$) can be converted to a 2-color binary tree isomorphism problem (2-BTI, comparing whether two trees are exactly isomorphic or not) under $AC^{0}$ reduction. Therefore, 2-color binary tree isomorphism is $NC^{1}$ hard. Because any solver of 2-BSI can solve 2-color binary tree isomorphism by simply adding a layer to check whether the two trees have same size, 2-BSI is harder than 2-BSI. Since $TC^{0}\subseteq NC^{1}$ and we assume that $TC^{0}\not=NC^{1}$, then for any fixed-length polynomial-parameter log-precision transformer, there exist a size of the 2-BSI problem such that the transformer can not handle. ∎
4.2 Divide-and-Conquer program guide Fixed-Depth Transformer to solve 2-BSI better than IO-prompting and Sequential Prompting
Theorem 4.2.
With Divide-and-Conquer prompting program, a log-precision transformer with fixed depth $L$ and fixed hidden dimension $d$ can solve the 2-BSI of any size with fixed-length prompt $m$ (for merge), $t$ (for sub-task tackling) and $d$ (for task decomposition). The context length and deconding length for each calling is $O(|t_{b}|)$ and the DaC prompting program will call the transformer $O(|t_{p}|)$ times.
Proof Sketch: The detailed proof to this theorem is provided in the Appendix A.1. Here we give a brief flow of the proof. To prove this theorem, we first show that there exists an problem size function $f(\cdot)$, hyper-parameter $w$, merge function $m(\cdot)$, sub-task tackling function $t(\cdot)$ and decomposition function $d(\cdot)$ that can solve the problem under the guidance of a divide-and-conquer program. The input length and output length of every aforementioned function are both $O(|t_{b}|)$, and the divide-and-conquer program will call them $O(|t_{p}|)$ times in total. Then we will prove that there exists one log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can express $a(\cdot)$, $t(\cdot)$ and $d(\cdot)$ respectively with different but fixed-length prompts. In this way, we can prove the theorem.
The above theorem explains why DaC program achieves better expressive power compared to IO-prompting. Meanwhile, from the above theoretic results, we can acquire a corollary:
Corollary 4.3.
A sequential prompting strategy (e.g., CoT and LtM) can simulate the trace of DaC prompting strategy to solve the 2-BSI problem. However, their average context length during decoding is $O(|t_{p}|\times|t_{b}|)$.
The above corollary reveals that when applying sequential prompting strategy such CoT and LtM to solve the problems with same algorithms like DaC, although both sequential prompting strategy and DaC prompting share the same decoding length, the maximum context length they require are different. Since DaC program follows the disentangled-sub-process principle, a transformer with context length of $O(t_{b})$ is sufficient. But for sequential prompting strategies, their maximum context length during decoding will be $O(|t_{p}|\times|t_{b}|)$. As claimed by (Yang & Ettinger, 2023), long context length during decoding will lead to the propagation of errors and hallucinations in the reasoning path.
@@ 5 Experiments
We evaluate the capacity of our prompting strategies on three kinds of tasks that are troubled by missing intermediate steps due to long inputs, even for state-of-the-art Language Models (e.g. GPT-3.5 and GPT-4): Multiplication and Addition of Large Numbers, Hallucination Detection in Long Context, and Fact-Checking for Misinformation Detection (Yang et al., 2023; Li et al., 2023a; Wadden et al., 2020). With these experiments, we present how Divide-and-Conquer (DaC) prompting handles this task better and suggest DaC could be a promising new paradigm in handling long documents or large inputs.
5.1 Multiplication of Long Integers
General-purpose Large Language Models, e.g. ChatGPT-3.5 and ChatGPT-4, have been bothered by its poor performance on arithmetic of big numbers, even though their expressive capacity has been proved to be enough for accurately calculate large number multiplication (Yang et al., 2023). In this experiment, we show that under zero-shot long integer multiplication, DaC can outperform other baseline prompting strategies.
Figure 3: Accuracy of our proposed method and different baseline prompting strategies on GPT-3.5 and GPT-4.
Task Setup: For this task, we randomly generated 200 pairs of 5-digit integers. We choose 5 for the digit length because according to previous works, ChatGPT-3.5 gets 0% accuracy on 4-digit multiplications (Cheng & Zhang, 2023), and ChatGPT-4 gets close to 0% accuracy on 5-digit multiplications (Yang et al., 2023). We evaluate the performance of calculation with two metrics: Accuracy and Edit Distance (Marzal & Vidal, 1993; Schaeffer et al., 2023). For accuracy, we count the result as correct one only every digit of it matches with the ground-truth answer. For edit distance, we calculate the minimum number of operations required to transform the output to the ground-truth.
Figure 4: Edit distances of DaC and other strategies on GPT-3.5 and GPT-4. Lower edit distance indicate better performance.
@@ 5.1 Multiplication of Long Integers
Table 1: Performance of different prompting methods on HaluEval dataset. We report the F1 score, Accuracy, Precision and Recall.
Setup of baselines and our method: In this task, our baselines include IO prompting, Chain of Thought (CoT), CoT-SC, Least-to-Most (LtM), and Decomposed Prompting (DeP). We also include our method’s variant: DaC without Disentangled-sub-process Principle (DaC w/o DP) for ablation study. Tree-of-Thoughts is not applicable. This is because that multiplication is deterministic calculation without requiring search in a tree. For our method, we apply Multi-Level Divide-and-Conquer program-guided solver. We set $f(a,b)$ as the minimum length of the two integers $a,b$. We set $w=2$ so that the recursion returns when both integers are no longer than $2$. If both integers are shorter, then the algorithm return. The prompts for decomposition asks the model to separate the two $n$-digit numbers to four $\frac{n}{2}$-digit numbers (e.g., separate $a,b$ as $a_{1},a_{2},b_{1},b_{2}$). The sub-task tackling prompt asks the model to calculate the multiplication of the four pairs of $\frac{n}{2}$-digit numbers and solution merge prompt merge them with bit shift and addition.
Results: Experimental results are shown in Fig. 3 and 4. As we can see, under all settings, our proposed prompting strategy outperform all the baselines. This indicate the advantage of our proposed method. Specifically, the thought-based methods, i.e. IO-Prompting, Chain-of-Thoughts and CoT-SC, all get accuracy close to 0 for both GPT-4 and GPT-3.5. The Least-to-Most prompting acquires better results on accuracy for GPT-3.5 and GPT-4 due to its advantage in disentangling task decomposition and task resolution. However, compared to our methods that can achieve accuracy better than 40% for both GPT-3.5 and GPT-4, these baselines significantly perform bad. For edit distance metric, the results are similar. For both GPT-3.5 and GPT-4, our method’s distances are lower than half of the baselines. Although for GPT-4, the Least-to-Most and DeP get edit distances that are significantly better than other baselines, it is still not comparable to our method. We claim that the advantage is brought by our method’s property in disentangling the task decomposition, task resolution and resolution assembly. The thought-based methods like CoT and CoT-SC intertwined the task decomposition and task resolution. As a result, its sub-task generation process is prone to be interrupted and gets wrong sub-task. Least-of-Most disentangle the the task decomposition and task resolution, and thus acquires much better results. However, its task resolution and resolution assembly processes are still intertwined and make the model prone to mistakes when assemblying the sub-task results. Also, DaC outperforms DaC w/o DP, indicating the importance of disentangled-sub-process principle.
5.2 Hallucination Detection in Long Context
@@
Task Setup: We use the HaluEval-Summary dataset. It is one of the three datasets in HaluEval benchmark for hallucination detection, which contains the hallucination generated by ChatGPT-3.5. HaluEval-Summary have the longest context and generated contents among all three tasks in this benchmark (Li et al., 2023a). Thus, detecting hallucination on this dataset requires repeatedly verify each sentence in the response, making standard prompting strategies acquire the worst accuracy across all three tasks. We report the Accuracy, F1 score [^2] We consider the hallucination pairs as positive samples., Precision score and Recall score.
Setup of baselines, ablation variants and our method: In this task, our baselines include IO prompting, Chain of Thought, CoT-SC, Tree-of-Thoughts Least-to-Most, and Decomposed Prompting. We also include our method’s variant: DaC without Disentangled-sub-process Principle (DaC w/o DP) for ablation study. In this task, the sub-tasks are verifying fragments of the summary, which are homogeneous and do not require recurssion. In such a setting, Decomposed Prompting is equivalent to LtM. For this task, we apply single level Divide-and-Conquer solver to decompose the summary to multiple sentences, handle them separately and then merge the conclusions of all sentences. The detailed prompts are provided in the Appendix.
Results: Experimental results are shown in Tab. 1. As we can see for all other baselines, the performance on GPT-3.5 is substantially worse than GPT-4, as the hallucinations are generated by GPT-3.5 and thus more deceptive to GPT-3.5. However, for our method, when we replace GPT-4 with GPT-3.5, although the performance still decrease, the drop is significantly lower than all other baselines, indicating the robustness of our proposed method toward deceptive contents. Also, for both GPT-3.5 and GPT-4, our proposed prompting strategy outperform the baselines, presenting the advantage of our method. More specifically, compared to IO-prompting, our model achieved better performance in general, indicating the advantage brought by stronger expressive power. Meanwhile, compared to CoT and CoT-SC results, our method clearly achieved much better recall. Tree-of-Thoughts, benefited by its searching ability, acquired significantly better recall score compared to other baselines. However, its significantly lower precision substantially harm its overall performance and leads to accuracy that is even worse than standard IO-prompting. Least-to-Most, which explicitly decomposing the task and enumerating all sub-tasks, achieves better recall compared to CoT and CoT-SC. However, it follows the flow of the hallucination, leading to proneness worse recall compared to Tree-of-Thoughts and our method. In contrary, our method carefully checked all sentences, locate the one containing factual error and merge the answers. Thus, our method balance the overall performance (F1 and Accuracy score). In addition, as we can see, although DaC w/o DP outperforms other baselines in general, it still perform worse than DaC. This phenomenon indicates the importance of disentangled-sub-process principles. Meanwhile, it is noticeable that the DaC w/o DP can be regarded as the upgrade version of LtM/DeP that replace sequential sub-tasks with parallel sub-tasks. Thus, the experimental results suggest that parallel sub-task strategy makes LLM more discerning for hallucination detection.
5.3 Fact-Verification for Misinformation Detection
The increasing abuse of misinformation toward manipulating public opinions on social media has been observed in different areas, such as healthcare (e.g. the recent COVID-19 pandemic) (Sharma et al., 2020, 2022). This threat is increasingly serious due to LLM’s capacity in content generation (Li et al., 2023b; Weidinger et al., 2021; Zhang et al., 2022). This challenge raise the importance of fact-verification, which aims at judging the authenticity of an article based on a collection of evidence from verified source (Whitehouse et al., 2022; Zhang & Gao, 2023). In this experiment, we present that DaC can outperform other baselines in fact-verification involved with news article.
@@ 5.3 Fact-Verification for Misinformation Detection
F1 G-Mean Prec Recall F1 G-Mean Prec Recall
@@
Table 2: Performance of different prompting methods on SciFact dataset. We report the F1 score, G-Mean score, Precision and Recall.
@@
Task Setup: In this experiment, we mainly adopt SciFact dataset (Wadden et al., 2020). In SciFact dataset, each sample is a pair of news and evidence, where the evidence is the abstract of a peer-reviewed paper and the news is a claim summarized by human annotators from fake news or true news. To better simulate the real-world scenario where news usually appears as an article, following Chen & Shu, we generate a dataset of article-level misinformation based on SciFact dataset. Specifically, for a given claim, we apply Large Language Model (i.e., ChatGPT-4) to extend the claim as an article based on the evidence. For this task, similar as hallucination detection, we apply single level Divide-and-Conquer solver to decompose the news article to multiple sentences, handle them separately and then merge the conclusions of all sentences. Also, the baselines in this experiments are the same as Hallucination Detection, for the same reason, Decomposed Prompting is equivalent to LtM. The evaluation metrics includes F1 score, G-Mean score (geometric mean of precision and recall), Precision score and Recall score. We do not apply accuracy as the positive and negative classes are not balanced.
@@
Results: Experimental results are shown in Tab. 2. Notably, GPT-3.5 incorporated with our proposed prompting strategy even outperform the performance of GPT-4 incorporated with IO-prompting, Least-to-Most, CoT and CoT-SC, which have significantly lower recall scores, indicating their proneness to deception. Only Tree-of-Thoughts, which is benefited by its advantage in exploring various options, acquired the best results among all baselines, but is still defeated by our proposed method. Moreover, as we can see, for GPT-4 the performance of CoT-SC is even worse than CoT, which is supposed to be a specific case of CoT-SC without exploration. These results suggests that, when facing deceptive contents generated on purpose, existing incremental works’ improvement may not be so robust.
6 Discussions and Limitations

### v6

@@ 4 Main Theoretic Results
2-color Binary Subtree Isomorphism problem is that, given a pattern 2-color binary tree $t_{p}$ and a base 2-color binary tree $t_{b}$, a solver is required to judge whether the pattern tree is isomorphic to a sub-tree of $t_{b}$
@@
There exists a log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can solve the 2-BSI of any size with fixed-length prompt $m$ (for merge), $t$ (for sub-task tackling) and $d$ (for task decomposition).
Proof Sketch: The detailed proof is provided in the Appendix A.2. Here we give a brief flow of the proof. To prove this theorem, we first show an algorithm that can solve the problem with divide-and-conquer strategy. Then we prove that there exists a log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can express the modules in the algorithms with different but fixed-length prompts. In this way, we can prove the theorem.
With the above theorem, we can prove that ${\sf NC^{1}}\subseteq S(DaC)$, which finishes the proof. With this theoretic results, we can answer the RQ 1:
Compared to IO prompting, DaC have theoretically guaranteed advantages in expressive power.
4.2 DaC vs. CoT
In this section, we compare Divide-and-Conquer with Chain-of-Thoughts in order to understand the utility and limitation of DaC prompting. The limitation of DaC prompting is that its expressive power is a subset of CoT prompting:
Proposition 4.3.
We denote the set of problems that a fixed-precision transformer with DaC prompting can tackle as $S(DaC)$. Similarly, we denote the set of problems that a fixed-precision transformer with CoT prompting can tackle as $S(CoT)$ Then we have the following results:
$S(DaC)\subseteq S(CoT)$ (2)
@@
The proof of this proposition is very straightforward. For any problem that DaC can solve, we can concatenate all outputs of LLM in dividing, tackling and merging as a sequence. Then we can prompt LLM with CoT to output this sequence. Thus, the problem set that DaC can resolve is a subset of CoT. ∎
The limitation revealed by the above theorem shows that compared to CoT, the appliance scope of Divide-and-Conquer is limited. However, by analyzing the average decoding context window size, we show that on specific tasks, divide and conquer can reduce the problem complexity:
Definition 2.
Decoding Context Window Size: In auto-regressive decoding, each token is decoded from a window that covers all previous tokens. We denote the length of the window as the Decoding Context Window Size of the token.
Proposition 4.4.
Suppose that a task contains $k$ sub-tasks, each of which does not rely on the answers of other sub-tasks. We define such sub-tasks as parallel sub-tasks. If an LLM tackle these sub-tasks sequentially with CoT, then the average decoding context window size of the sub-tasks’ resolution will be $C+\frac{\sum_{i=1}^{k}r_{i}-1}{2}$, where $r_{i}$ is the length of the response to the $i$-th sub-task and $C$ is the length of input context. If we tackle them parallely with DaC, then the average decoding context window size of the sub-tasks’ resolution will be $C+\sum_{i=1}^{k}\frac{(r_{i}-1)^{2}}{2\sum_{j=1}^{k}r_{j}}<C+\frac{\sum_{i=1}^{k}r_{i}-1}{2}$.
The above proposition shows that when task contains a large amount of parallel sub-tasks, DaC is more helpful for reducing the average decoding context window size than CoT. Existing works have empirically showed that long decoding context window will propagate intermediate errors and thus increase the probability of generating hallucination (Yang and Ettinger, 2023). Thus, we can acquire a conclusion that DaC is competetive on tasks that contain a large amount of parallel sub-tasks and are bothered by intermediate errors and hallucination. With these theoretic results, we can answer the RQ 2:
Compared to CoT and its variants, DaC prompting’s expressive power is weaker. However, on tasks containing a large amount of parallel sub-tasks, DaC is more helpful.
4.3 Advantages of DaC
The above analysis answer the two research questions that we proposed. By summarizing these two answers, we can acquire the two conditions such that when a task simultaneously satisfied both conditions, DaC bring performance boost:
  • Condition 1: the task is harder than $S(IO)$, such as ${\sf TC^{0}}$-complete problems and ${\sf NC^{1}}$-complete problems.
  • Condition 2: the task contains a large amount of parallel sub-tasks and is bothered by hallucinations or intermediate errors.
In Tab. 1, we present some sample tasks that satisfied the conditions. Also, we list some tasks that typically do not satisfy the conditions. This is helpful for guiding prompt engineering. Details are provided in Appendix A.6.
Applicable Tasks Non-Applicable Tasks
Integer Multiplication Integer Addition
Fact Verification Multi-round QA
Consistency Evaluation Planning
Table 1: We list some exaple tasks that satisfy the conditions and some tasks that do not satisfy the conditions.
@@ 5 Experiments
5.1 Case 1: Long Integer Arithmetic
In this case, we consider two tasks in long integer arithmetic: Multiplication, which satisfy the conditions we proposed, and Addition, which does not satisfy the first condition  [^1] Multiplication is a ${\sf TC^{0}}$-complete problem and can be divided to multiple parallel sub-tasks, while Addition is in $S(IO)$Barcelo et al. (2023). Our experiment results will show that DaC prompting bring performance boost on multiplication and does not bring boost on integer addition.
Figure 3: Edit distance of DaC and baseline prompting strategies on GPT-3.5 and GPT-4 for Multiplication.
Figure 4: Edit distance of DaC and baseline prompting strategies on GPT-3.5 and GPT-4 for Addition.
@@ 5.1 Case 1: Long Integer Arithmetic
Table 2: Performance of different prompting methods on HaluEval dataset.
Setup of baselines and DaC: In this task, our baselines include IO prompting, Chain of Thought (CoT), CoT-SC, Least-to-Most (LtM), and Decomposed Prompting (DeP). Tree-of-Thoughts is not applicable. This is because that multiplication is deterministic calculation without requiring search in a tree. For DaC, we apply Multi-Level Divide-and-Conquer program-guided solver.
Results: Experimental results are shown in Fig. 3 and 4. As we can see, for integer addition which does not satisfy our proposed conditions, the performance of DaC, CoT and its variants does not significantly outperform IO prompting for both ChatGPT-3.5 and 4. However, for integer multiplication which satisfy our proposed conditions, under all settings, our proposed prompting strategy outperform all the baselines. This phenomenon indicate that our proposed conditions are useful for recognizing the tasks where DaC is more powerful.
5.2 Case 2: Fact Verification of Long Text
In the previous section, we show that for arithmetic tasks, our proposed conditions are discerning to the tasks where divide-and-conquer has advantages. In this section, we further present our conditions can be applied to natural language tasks. Specifically, we present the performance of baselines and Divide-and-Conquer on fact verification of long text. In this task, the LLM is required to whether a long corpus is aligned with base knowledge. This task satisfied the proposed two conditions. For the first condition, we can reduce a 2-BTI problem to fact verification by describing the two trees with natural language. In this way, we can convert the trees to two paragraphs and what we need to do is to ask the LLM to judge whether the two paragraphs are aligned or not. For the second condition, since we are tackling long text, then each sentence can be regarded as parallel sub-tasks. We select two benchmarks of fact verification: Fact-Verification for Hallucination Detection and Fact-Verification for Misinformation Detection
5.2.1 Hallucination Detection
@@
F1 G-M Prec Recall F1 G-M Prec Recall
@@
Table 3: Performance of different prompting methods on SciFact dataset.
Task Setup: We use the HaluEval-Summary dataset. It is one of the three datasets in HaluEval benchmark for hallucination detection, which contains the hallucination generated by ChatGPT-3.5. HaluEval-Summary have the longest context and generated contents among all three tasks in this benchmark (Li et al., 2023a). Thus, detecting hallucination on this dataset requires repeatedly verify each sentence in the response, making standard prompting strategies acquire the worst accuracy across all three tasks. We report the Accuracy, F1 score (the hallucination pairs are positive samples), Precision and Recall.
Setup of baselines, ablation variants and DaC: In this task, our baselines include IO prompting, Chain of Thought, CoT-SC, Tree-of-Thoughts Least-to-Most, and Decomposed Prompting. In this task, the sub-tasks are verifying fragments of the summary, which are homogeneous and do not require recurssion. In such a setting, Decomposed Prompting is equivalent to LtM. For this task, we apply single level Divide-and-Conquer solver to decompose the summary to multiple sentences, handle them separately and then merge the conclusions of all sentences. The details are in Appendix.
Results: Experimental results are shown in Tab. 2. For both GPT-3.5 and GPT-4, our proposed prompting strategy outperform the baselines, presenting the advantage of DaC. More specifically, compared to IO-prompting, DaC achieved better performance in general, indicating the advantage brought by stronger expressive power. Meanwhile, compared to CoT and CoT-SC results, DaC clearly achieved much better recall. Tree-of-Thoughts, benefited by its searching ability, acquired significantly better recall score compared to other baselines. However, its significantly lower precision substantially harm its overall performance and leads to accuracy that is even worse than standard IO-prompting. In contrary, DaC carefully checked all sentences, locate the one containing factual error and merge the answers.
5.2.2 Misinformation Detection
The increasing abuse of misinformation toward manipulating public opinions on social media has been observed in different areas, such as healthcare (e.g. the recent COVID-19 pandemic) (Sharma et al., 2020, 2022). This threat is increasingly serious due to LLM’s capacity in content generation (Li et al., 2023b; Weidinger et al., 2021; Zhang et al., 2022). This challenge raise the importance of fact-verification, which aims at judging the authenticity of an article based on a collection of evidence from verified source (Whitehouse et al., 2022; Zhang and Gao, 2023). In this experiment, we present that DaC can outperform other baselines in fact-verification involved with news article.
@@ 5.2.2 Misinformation Detection
Task Setup: In this experiment, we mainly adopt SciFact dataset (Wadden et al., 2020). In SciFact dataset, each sample is a pair of news and evidence, where the evidence is the abstract of a peer-reviewed paper and the news is a sentence of claim. To better simulate the real-world scenario where news on social media usually appears as an paragraph of post, following Chen and Shu, we generate a dataset of paragraph-level misinformation based on SciFact dataset. Specifically, for a given claim, we apply ChatGPT-4 to extend the claim as an article based on the evidence. For this task, similar as hallucination detection, we apply single level Divide-and-Conquer solver to decompose the news article to multiple sentences, handle them separately and then merge the conclusions of all sentences. Also, the baselines in this experiments are the same as Hallucination Detection. The evaluation metrics includes F1 score, G-Mean score (geometric mean of precision and recall), Precision and Recall. We do not apply accuracy as the positive and negative classes are not balanced.
@@
Results: Experimental results are shown in Tab. 3. Notably, GPT-3.5 incorporated with our proposed prompting strategy even outperform the performance of GPT-4 incorporated with IO-prompting, Least-to-Most, CoT and CoT-SC, which have significantly lower recall scores, indicating their proneness to deception. Only Tree-of-Thoughts, which is benefited by its advantage in exploring various options, acquired the best results among all baselines, but is still defeated by DaC. Moreover, as we can see, for GPT-4 the performance of CoT-SC is even worse than CoT, which is supposed to be a specific case of CoT-SC without exploration. These results suggests that, when facing deceptive contents generated on purpose, existing works’ improvement may not be robust.
6 Conclusions
In this paper, we analyze the utility and limitations of divide-and-conquer prompting strategy. We first provide theoretic analysis to Divide-and-Conquer prompting and compare it with representative prompting strategies. Based on these theoretic results, we summarize two conditions under which a task is suitable for Divide-and-Conquer prompting. After that we present empirical results that validated our theoretic analysis.
Appendix A Appendix
A.1 Discussions and Limitations